OpenStudy (anonymous):

Find the x-coordinates where f '(x) = 0 for f(x) = 2x – sin(2x) in the interval [0, 2π].

OpenStudy (anonymous):

@Luigi0210 can u plz help on this 1? :) its not a word problem

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

take the derivative set it equal to zero and solve it is not a word problem

OpenStudy (anonymous):

okay so i take the derivative of 2x-sin(2x), can u plz show me those steps? i like to see it all written out it helps me understand better.

OpenStudy (anonymous):

do you know what the derivative of \(2x\) is? `

OpenStudy (anonymous):

yes 2

OpenStudy (anonymous):

Idk the derivative of -sin(2x)

OpenStudy (anonymous):

ok that was the first step how about the derivative of \(-\sin(2x)\)?

OpenStudy (anonymous):

-2cos(2x)

OpenStudy (anonymous):

?:)

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@welshfella @pooja195 @Luigi0210 plz help!!!!!! :(

OpenStudy (anonymous):

yes you have it

OpenStudy (anonymous):

solve \[2-2\cos(2x)=0\] and you are done

OpenStudy (anonymous):

What do you mean solve? what about the interval [0,2pi]?

OpenStudy (anonymous):

what do you mean "what do you mean" solve for \(x\) there is no other way i can think of to say it

OpenStudy (anonymous):

start with \[2=2\cos(2x)\] so \[\cos(2x)=1\]

OpenStudy (anonymous):

then solve that one more or less in your head for the first value

OpenStudy (anonymous):

do we divide both sides by cos?

OpenStudy (anonymous):

so it'd be 1/cos = 2x or can we not separate them

OpenStudy (anonymous):

to get \[2x=\frac{1}{\text{cos}}\]?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

a bit confused right?

OpenStudy (anonymous):

yes just at the solving for x part

OpenStudy (anonymous):

\[f(x)=x+2\\ x=\frac{x+2}{f}\]

OpenStudy (anonymous):

so i have: cos(2x)=1

OpenStudy (anonymous):

cosine is a function you are trying to solve \[\cos(2x)=1\] think of a number whose cosine is 1

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and if \(2x=0\) then \(x=0\) now we get another one

OpenStudy (anonymous):

\[\cos(2\pi)=1\] also, so if \[2x=2\pi\] then \[x=\pi\]

OpenStudy (anonymous):

those are the two in your interval from zero to two pi

OpenStudy (anonymous):

ohh so the coordinates are 0 & pi?

OpenStudy (anonymous):

the first coordinates are, yes

OpenStudy (anonymous):

lol there's more???????

OpenStudy (anonymous):

@perl @perl @perl can u plz help ? :)